# Eventually LIL Regret: Almost Sure $\ln\ln T$ Regret for a sub-Gaussian Mixture on Unbounded Data

### v1

@@ 1 Introduction
    If in Point 2 above, the data are assumed to be drawn from a distribution (not necessarily iid) such that the mixture processes is a non-negative supermartingale (Definition 1), the “Ville event” $\mathcal{E}_{\alpha}$ has measure at least $1-\alpha$. Thus, the pathwise regret bound from Point 2 holds for a large fraction of paths (Theorem 4.1). Importantly, the stochasticity is only used to quantify the size of $\mathcal{E}_{\alpha}$ — the regret bounds still hold deterministically on $\mathcal{E}_{\alpha}$. Our deterministic bounds thus imply the high probability regret bound: $P(R_{T}=O(\ln\ln V_{T}+\ln^{2}(1/\alpha)/V_{T}+\log(1/\alpha))\geq 1-\alpha$ simultaneously for a large class of measures $P$ that ensure $Z_{t}$ is a sub-Gaussian process.
@@ 2.1 Setup
$Z_{t}:=\int_{\mathbb{R}}e^{f_{t}(\eta)}\pi(\eta)d\eta,$
with $Z_{0}=1$; and $Z_{t}:=1$ for $t$ such that $V_{t}=0$. The regret of $Z_{t}$ with respect to the maximum attainable wealth is defined as
$R_{t}:=L_{t}^{*}-\ln Z_{t}.$
@@
$\mathcal{E}_{\alpha}=\left\{\sup\limits_{t\geq 1}~\ln Z_{t}\leq\ln\frac{1}{\alpha}\right\},$
on which the log-wealth $Z_{t}$ remains bounded. We call $\mathcal{E}_{\alpha}$ the “Ville event”.
@@
The main focus of this work is to establish regret bounds for the mixture wealth process $Z_{t}$, constructed using a mixing distribution $\pi$ originally introduced in Robbins and Siegmund [1968]. We show that this classical stochastic construction can be reinterpreted within an adversarial framework: specifically, $Z_{t}$ admits deterministic, path-wise regret guarantees, which have an interesting nontrivial interpretation. The worst-case regret is unbounded (linear in $t$ or $V_{t}$); but we can prove that one attains low regret on every path in $\mathcal{E}_{\alpha}$. What makes $\mathcal{E}_{\alpha}$ special is that under natural stochastic assumptions (implied by the underlying mixture processes), the event holds with high probability. So we obtain a deterministic low regret bound on every path in a high probability event. In fact, we even show that we achieve a small regret eventually on almost every path (i.e. on every path in measure $1$ set $\mathcal{E}_{0}$). In other words, the set of paths on which regret is unbounded is a measure zero set (under many natural families of distributions specified later).
@@ 2.3 Literature Review
Cutkosky and Sarlos [2019] use wealth/betting arguments to derive a path-wise regret bound in an adversarial online learning setting. However, they require that the gradients are bounded in a certain dual norm.
Orabona and Jun [2023] leverage the wealth-regret connection, but their primary aim is to construct time-uniform confidence sets for the mean of bounded data using the wealth process of certain betting strategies. In their work, they implicitly prove iterated logarithmic conditional regret bounds for bounded data for a wealth process different from $Z_{t}$. By contrast, we extend the analysis to unbounded data and also prove regret bounds for $Z_{t}$ that hold for arbitrary sequences of observations.
@@ 3 Warm-up: Logarithmic Conditional Regret
$Z_{t}=\frac{1}{\sqrt{1+\sigma_{0}^{2}V_{t}}}e^{\frac{\sigma_{0}^{2}S_{t}^{2}}{2(1+\sigma_{0}^{2}V_{t})}},$
and as earlier, $Z_{t}=1$ for $t$ such that $V_{t}=0$. Theorem 3.1 below shows that the mixture $Z_{t}$ incurs a regret of $O(\ln V_{t}+\ln(1/\alpha))$ on the set $\mathcal{E}_{\alpha}$, defined in (2.1).
@@
The proof of the above theorem proceeds by first establishing an expression for the unconditional regret, which corresponds to the first expression in the theorem statement. Next, on the set $\mathcal{E}_{\alpha}$, $\ln Z_{t}$ is bounded. Using this bound in the expression for $R_{t}$ after appropriate rearrangements, we get a conditional bound on $R_{t}$ on the set $\mathcal{E}_{\alpha}$, in terms of $V_{t}$ and $S_{t}$. The expression in (3.1) then follows by bounding certain occurrences of $V_{t}$ by $v_{0}$. The proof is straightforward, and we present it in Appendix B for completeness.
@@
$\mathcal{E}_{0}:=\left\{V_{t}\uparrow\infty;~\limsup_{t}Z_{t}<\infty\right\}.$
@@
i.e. on every path in $\mathcal{E}_{0}$, and for every constant $d$, there exists an $n$ such that for all $t\geq n$, $R_{t}\leq\ln(1+V_{t})+d$. In particular, dividing by $V_{t}$ and taking the limit as $t\to\infty$, we get that $S_{t}/V_{t}\to 0$ path-wise on $\mathcal{E}_{0}$. Finally, if the data are drawn from a distribution $P$ such that $(S_{t})_{t\geq 1}$ is sub-Gaussian with variance proxy $(V_{t})_{t\geq 1}$, then $Z_{t}$ is a nonnegative super-martingale, immediately implying that $\mathcal{E}_{0}$ is a measure 1 event, thus implying the (self-normalized) SLLN. Contrapositively, if there is a path on which $S_{t}/V_{t}\to 0$ does not hold, then $\limsup_{t}Z_{t}=\infty$ on that path.
Ville’s martingale theorem [Ville, 1939] states that fixing a measure $P$, for every event $A$ such that $P(A)=0$, there exists a nonnegative $P$-martingale $M^{P}$ that increases to $\infty$ on $A$. But this $M^{P}$ is potentially different for every $P$. Here, we have a single super-martingale $Z=(Z_{t})_{t\geq 1}$ that serves as a witness for the SLLN for an entire class of distributions $\mathcal{P}$ (eg, zero mean, 1-sub-Gaussian distributions). In general, martingales and even super-martingales may not suffice as witnesses for the strong law for composite $\mathcal{P}$, or more generally for a composite generalization of Ville’s theorem; for example, Ruf et al. [2023] showed that in general one must employ e-processes instead (see Ruf et al. [2023] for a definition). A general characterization of when super-martingales suffice is open; the above result provides an interesting example.
Moreover, Theorem 3.3 provides an explicit witness super-martingale of the violation of the strong law, meaning that if the strong law is violated on some path, $\limsup_{t}Z_{t}=\infty$ on that path.
@@
$Z_{t}:=\int_{\mathbb{R}}e^{f_{t}(\eta)}\pi(\eta)d\eta=\frac{e^{\frac{S_{t}^{2}}{2(1+V_{t})}}}{\sqrt{1+V_{t}}},$
@@ 4 Main Result: Iterated Logarithmic Conditional Regret
Recall $S_{t}$, $V_{t}$, $\eta^{*}_{t}$, $L^{*}_{t}$, and $\mathcal{E}_{\alpha}$ from previous sections. In this section, we establish a regret bound for Robbins’ mixture strategy [Robbins and Siegmund, 1968], conditioned on the set $\mathcal{E}_{\alpha}$ from (2.1) on which the log of the mixture wealth remains bounded. To achieve this, as in Section 3, we first prove a path-wise regret bound that holds for all data sequences $X_{1},X_{2},\dots$. In general, this bound can be arbitrarily large (at least linear in $t$). However, we show that on the set $\mathcal{E}_{\alpha}$, where the mixture process $Z_{t}$ is restricted to be uniformly bounded, the regret is also small. In fact, for the mixture wealth using Robbins’ heavy-near-zero prior, we show that the bound on regret till time $t$, for every $t$, is $O(\ln\ln V_{t})$.
To formalize, in this section, we consider the mixture $\pi$, which, for $c\geq 6.6e$ and $\eta\in[-1,1]$, is given by
$\pi(\eta)=\frac{1}{Z_{0}}\frac{{\mathbf{1}}_{|\eta|\leq 1}}{|\eta|(\ln\tfrac{c}{|\eta|})(\ln\ln\tfrac{c}{|\eta|})^{2}},$
and equals 0 otherwise. Here, $Z_{0}$ is the normalization constant
@@
Note that in the above discussion, we assumed the data to be stochastic only to develop some intuition about the bound in (5). However, it does not require any stochasticity assumption, and instead holds path-wise. Next, in (4.1) we show that on the set where the data is well-behaved in the sense that the wealth process $Z_{t}$ is uniformly bounded from above, the path wise regret is small, and is in fact, $O(\ln\ln V_{t})$ at time $t$. Even the path-wise bound in (4.1) does not require any stochasticity assumption. Stochasticity is only used in the last part of the theorem, which shows that if the data is stochastic and satisfies certain conditions, the set $\mathcal{E}_{\alpha}$ where (4.1) holds is indeed a large set (of probability at least $1-\alpha$).
@@
Next, under the stochasticity assumption, we observe that our regret bound in (5) immediately implies a $(1-\alpha)$-CS (a sequence of confidence intervals with a time-uniform coverage guarantee of $1-\alpha$) for the mean of the unknown data-generating distribution. To see this, let $B_{t}$ denote the right hand side of (5), and recall that $R_{t}={S^{2}_{t}}/({2V_{t}})-\log Z_{t}$. Then, from (5), the following holds path wise on $\mathcal{E}_{\alpha}$:
@@
Theorem 4.1 above proves a path-wise regret bound (Eq. (5)), as well as a conditional regret bound of $O(\ln\ln V_{t})$ (Eq. (4.1)) that holds for every $t$ on a set $\mathcal{E}_{\alpha}$. If the data is stochastic, then $\mathcal{E}_{\alpha}$ is large. In the following theorem, we show that the $O(\ln\ln V_{t})$ conditional regret bound holds eventually on a larger set $\mathcal{E}_{0}$ where the process $Z_{t}$ remains finite, but possibly not uniformly bounded. This further implies that on $\mathcal{E}_{0}$, $|S_{t}|/V_{t}\to 0$ and $\limsup_{t}|S_{t}|/\sqrt{V_{t}\ln\ln V_{t}}\leq\sqrt{2}$. In other words, either both of these aforementioned convergences hold or the process $Z_{t}$ explodes. In particular, if the data are stochastic with appropriate assumptions, the set $\mathcal{E}_{0}$ is a set of measure $1$, and we conclude that the process $Z_{t}$ acts as a witness for the self-normalized SLLN and (upper) LIL.
@@
$\mathcal{E}_{0}:=\left\{V_{t}\uparrow\infty;~\limsup_{t}Z_{t}<\infty\right\}.$
@@
path-wise on $\mathcal{E}_{0}$. As a corollary, if the data are drawn from a distribution such that $(S_{t})_{t\geq 1}$ is a sub-Gaussian process with variance $(V_{t})_{t\geq 1}$, then $Z_{t}$ is a nonnegative super-martingale, immediately implying that $\mathcal{E}_{0}$ is a measure 1 event, thus implying the self-normalized SLLN and LIL.
As in Theorem 3.3, Theorem 4.2 demonstrates a deterministic statement of the form: either $R_{t}=O(\ln\ln V_{t})$, or the wealth process $Z_{t}$ explodes. In case of stochastic data, it provides an explicit super-martingale witnessing the violation of the SLLN and LIL: if either of these is violated on some path, $\limsup_{t}Z_{t}=\infty$ on that path.
@@
Since $R_{t}=\frac{S^{2}_{t}}{2V_{t}}-\ln Z_{t}$, from (5) we get a lower bound on $Z_{t}$. Further, using $\ln(x)\leq x-1$ and completing the squares, we get the following path-wise lower bound on $Z_{t}$:
$\ln Z_{t}\geq\begin{cases}\frac{S_{t}^{2}}{2V_{t}}-\frac{1}{2}-\ln\tfrac{2}{\ln\ln c}-\ln\ln(c\sqrt{1+V_{t}})-2\ln\ln\ln(c\sqrt{1+V_{t}}),&\text{ if }\frac{|S_{t}|}{V_{t}}\leq\frac{1}{\sqrt{1+V_{t}}}\\ \frac{1}{2}\left(\!\frac{|S_{t}|}{\sqrt{V_{t}}}-\sqrt{1+\frac{1}{V_{t}}}\right)^{2}-\ln\tfrac{2}{\ln\ln c}-\frac{1}{2V_{t}}\\ \qquad\qquad-\ln\ln\left(c\sqrt{1+V_{t}}\right)-2\ln\ln\ln(c\sqrt{1+V_{t}}),&\text{ if }\frac{1}{\sqrt{1+V_{t}}}<\frac{|S_{t}|}{V_{t}}\leq 1\\ \frac{S_{t}^{2}}{2V_{t}}-\frac{V_{t}}{2}\left(\frac{|S_{t}|}{V_{t}}-1+\rho\right)^{2}+\ln\left(\frac{\rho\ln\ln c}{\ln({c}/{1-\rho})(\ln\ln({c}/{1-\rho}))^{2}}\right),&\text{ if }1<\frac{|S_{t}|}{V_{t}}.\end{cases}$ (11)
@@
$\ln Z_{t}\geq V_{t}\left(\frac{1-\rho^{2}}{2}\right)+\ln\left(\frac{\rho\ln\ln c}{\left(\ln\tfrac{c}{1-\rho}\right)\left(\ln\ln\tfrac{c}{1-\rho}\right)^{2}}\right).$
Recall that on $\mathcal{E}_{0}$, $V_{t}\uparrow\infty$ and $\limsup_{t}Z_{t}<\infty$. From this and the inequality above it, we see that on $\mathcal{E}_{0}$, if $\limsup_{t}\frac{|S_{t}|}{V_{t}}>1$ (i.e., the third branch occurs infinitely often), then $\limsup_{t}\ln Z_{t}=\infty$, which leads to a contradiction.
@@
Recall from Section 2.1 that $R_{t}:=f_{t}(\eta^{*}_{t})-\ln Z_{t}$, and that for $t$ such that $V_{t}=0$, we also have $S_{t}=0$, and hence for such $t$, for all $\eta\in\mathbb{R}$, $f_{t}(\eta)=0$, and $\ln Z_{t}=0$. Thus, the inequalities in the theorem hold trivially in this case. Henceforth, we focus on $t$ so that $V_{t}>0$.

### v2

@@ 1 Introduction
    If in Point 2 above, the data are assumed to be drawn from a distribution (not necessarily iid) such that the mixture processes is a non-negative supermartingale (Definition 1), the “Ville event” $\mathcal{E}_{\alpha}$ has measure at least $1-\alpha$. Thus, the pathwise regret bound from Point 2 holds for a large fraction of paths (Theorem 4.1). Importantly, the stochasticity is only used to quantify the size of $\mathcal{E}_{\alpha}$ — the regret bounds still hold deterministically on $\mathcal{E}_{\alpha}$. Our deterministic bounds thus imply the high probability regret bound: $P(R_{T}=O(\ln\ln V_{T}+\ln^{2}(1/\alpha)/V_{T}+\log(1/\alpha))\geq 1-\alpha$ simultaneously for a large class of measures $P$ that ensure $W_{t}$ is a sub-Gaussian process.
@@ 2.1 Setup
$W_{t}:=\int_{\mathbb{R}}e^{f_{t}(\eta)}\pi(\eta)d\eta,$
with $W_{0}=1$; and $W_{t}:=1$ for $t$ such that $V_{t}=0$. The regret of $W_{t}$ with respect to the maximum attainable wealth is defined as
$R_{t}:=L_{t}^{*}-\ln W_{t}.$
@@
$\mathcal{E}_{\alpha}=\left\{\sup\limits_{t\geq 1}~\ln W_{t}\leq\ln\frac{1}{\alpha}\right\},$
on which the log-wealth $W_{t}$ remains bounded. We call $\mathcal{E}_{\alpha}$ the “Ville event”.
@@
The main focus of this work is to establish regret bounds for the mixture wealth process $W_{t}$, constructed using a mixing distribution $\pi$ originally introduced in Robbins and Siegmund [1968]. We show that this classical stochastic construction can be reinterpreted within an adversarial framework: specifically, $W_{t}$ admits deterministic, path-wise regret guarantees, which have an interesting nontrivial interpretation. The worst-case regret is unbounded (linear in $t$ or $V_{t}$); but we can prove that one attains low regret on every path in $\mathcal{E}_{\alpha}$. What makes $\mathcal{E}_{\alpha}$ special is that under natural stochastic assumptions (implied by the underlying mixture processes), the event holds with high probability. So we obtain a deterministic low regret bound on every path in a high probability event. In fact, we even show that we achieve a small regret eventually on almost every path (i.e. on every path in measure $1$ set $\mathcal{E}_{0}$). In other words, the set of paths on which regret is unbounded is a measure zero set (under many natural families of distributions specified later).
@@ 2.3 Literature Review
Cutkosky and Sarlos [2019] use wealth/betting arguments to derive a path-wise regret bound in an adversarial online learning setting. However, they require that the gradients are bounded in a certain dual norm. Koolen and Van Erven [2015] prove a deterministic, pathwise regret bound that is $O(\sqrt{V_{t}\ln\ln V_{t}})$ for a certain variance process $V_{t}\leq t$. However, this result crucially relies on bounded losses.
Orabona and Jun [2023] leverage the wealth-regret connection, but their primary aim is to construct time-uniform confidence sets for the mean of bounded data using the wealth process of certain betting strategies. In their work, they implicitly prove iterated logarithmic conditional regret bounds for bounded data for a wealth process different from $W_{t}$. By contrast, we extend the analysis to unbounded data and also prove regret bounds for $W_{t}$ that hold for arbitrary sequences of observations.
@@ 3 Warm-up: Logarithmic Conditional Regret
$W_{t}=\frac{1}{\sqrt{1+\sigma_{0}^{2}V_{t}}}e^{\frac{\sigma_{0}^{2}S_{t}^{2}}{2(1+\sigma_{0}^{2}V_{t})}},$
and as earlier, $W_{t}=1$ for $t$ such that $V_{t}=0$. Theorem 3.1 below shows that the mixture $W_{t}$ incurs a regret of $O(\ln V_{t}+\ln(1/\alpha))$ on the set $\mathcal{E}_{\alpha}$, defined in (2.1).
@@
The proof of the above theorem proceeds by first establishing an expression for the unconditional regret, which corresponds to the first expression in the theorem statement. Next, on the set $\mathcal{E}_{\alpha}$, $\ln W_{t}$ is bounded. Using this bound in the expression for $R_{t}$ after appropriate rearrangements, we get a conditional bound on $R_{t}$ on the set $\mathcal{E}_{\alpha}$, in terms of $V_{t}$ and $S_{t}$. The expression in (3.1) then follows by bounding certain occurrences of $V_{t}$ by $v_{0}$. The proof is straightforward, and we present it in Appendix B for completeness.
@@
$\mathcal{E}_{0}:=\left\{V_{t}\uparrow\infty;~\limsup_{t}W_{t}<\infty\right\}.$
@@
i.e. on every path in $\mathcal{E}_{0}$, and for every constant $d$, there exists an $n$ such that for all $t\geq n$, $R_{t}\leq\ln(1+V_{t})+d$. In particular, dividing by $V_{t}$ and taking the limit as $t\to\infty$, we get that $S_{t}/V_{t}\to 0$ path-wise on $\mathcal{E}_{0}$. Finally, if the data are drawn from a distribution $P$ such that $(S_{t})_{t\geq 1}$ is sub-Gaussian with variance proxy $(V_{t})_{t\geq 1}$, then $W_{t}$ is a nonnegative super-martingale, immediately implying that $\mathcal{E}_{0}$ is a measure 1 event, thus implying the (self-normalized) SLLN. Contrapositively, if there is a path on which $S_{t}/V_{t}\to 0$ does not hold, then $\limsup_{t}W_{t}=\infty$ on that path.
Ville’s martingale theorem [Ville, 1939] states that fixing a measure $P$, for every event $A$ such that $P(A)=0$, there exists a nonnegative $P$-martingale $M^{P}$ that increases to $\infty$ on $A$. But this $M^{P}$ is potentially different for every $P$. Here, we have a single super-martingale $Z=(W_{t})_{t\geq 1}$ that serves as a witness for the SLLN for an entire class of distributions $\mathcal{P}$ (eg, zero mean, 1-sub-Gaussian distributions). In general, martingales and even super-martingales may not suffice as witnesses for the strong law for composite $\mathcal{P}$, or more generally for a composite generalization of Ville’s theorem; for example, Ruf et al. [2023] showed that in general one must employ e-processes instead (see Ruf et al. [2023] for a definition). A general characterization of when super-martingales suffice is open; the above result provides an interesting example.
Moreover, Theorem 3.3 provides an explicit witness super-martingale of the violation of the strong law, meaning that if the strong law is violated on some path, $\limsup_{t}W_{t}=\infty$ on that path.
@@
$W_{t}:=\int_{\mathbb{R}}e^{f_{t}(\eta)}\pi(\eta)d\eta=\frac{e^{\frac{S_{t}^{2}}{2(1+V_{t})}}}{\sqrt{1+V_{t}}},$
@@ 4 Main Result: Iterated Logarithmic Conditional Regret
Recall $S_{t}$, $V_{t}$, $f_{t}$, $\eta^{*}_{t}$, $L^{*}_{t}$, and $\mathcal{E}_{\alpha}$ from Section 2.1. In this section, we establish a regret bound for Robbins’ mixture strategy [Robbins and Siegmund, 1968], conditioned on the set $\mathcal{E}_{\alpha}$ from (2.1) on which the log of the mixture wealth remains bounded. To achieve this, as in Section 3, we first prove a path-wise regret bound that holds for all data sequences $X_{1},X_{2},\dots$. In general, this bound can be arbitrarily large (at least linear in $t$). However, we show that on the set $\mathcal{E}_{\alpha}$, where the mixture process $W_{t}$ is restricted to be uniformly bounded, the regret is also small. In fact, for the mixture wealth using Robbins’ heavy-near-zero prior, we show that the bound on regret till time $t$, for every $t$, is $O(\ln\ln V_{t})$.
To formalize, for $f_{t}$ defined in (2.1), in this section, we consider the mixture wealth process $W_{t}$ given by
$W_{t}=\frac{1}{Z_{0}}\int\limits_{-1}^{1}\frac{e^{f_{t}(\eta)}}{|\eta|(\ln\tfrac{c}{|\eta|})(\ln\ln\tfrac{c}{|\eta|})^{2}}d\eta,$
which is obtained by using the following mixture $\pi$ in (2.1): for $c\geq 6.6e$ and $\eta\in[-1,1]$,
$\pi(\eta)=\frac{1}{Z_{0}}\frac{{\mathbf{1}}_{|\eta|\leq 1}}{|\eta|(\ln\tfrac{c}{|\eta|})(\ln\ln\tfrac{c}{|\eta|})^{2}}.$
It equals 0 outside [1,1][-1,1]. Here, $Z_{0}$ is the normalization constant
@@
Note that in the above discussion, we assumed the data to be stochastic only to develop some intuition about the bound in (7). However, it does not require any stochasticity assumption, and instead holds path-wise. Next, in (4.1) we show that on the set where the data is well-behaved in the sense that the wealth process $W_{t}$ is uniformly bounded from above, the path wise regret is small, and is in fact, $O(\ln\ln V_{t})$ at time $t$. Even the path-wise bound in (4.1) does not require any stochasticity assumption. Stochasticity is only used in the last part of the theorem, which shows that if the data is stochastic and satisfies certain conditions, the set $\mathcal{E}_{\alpha}$ where (4.1) holds is indeed a large set (of probability at least $1-\alpha$).
@@
Next, under the stochasticity assumption, we observe that our regret bound in (7) immediately implies a $(1-\alpha)$-CS (a sequence of confidence intervals with a time-uniform coverage guarantee of $1-\alpha$) for the mean of the unknown data-generating distribution. To see this, let $B_{t}$ denote the right hand side of (7), and recall that $R_{t}={S^{2}_{t}}/({2V_{t}})-\log W_{t}$. Then, from (7), the following holds path wise on $\mathcal{E}_{\alpha}$:
@@
Theorem 4.1 above proves a path-wise regret bound (Eq. (7)), as well as a conditional regret bound of $O(\ln\ln V_{t})$ (Eq. (4.1)) that holds for every $t$ on a set $\mathcal{E}_{\alpha}$. If the data is stochastic, then $\mathcal{E}_{\alpha}$ is large. In the following theorem, we show that the $O(\ln\ln V_{t})$ conditional regret bound holds eventually on a larger set $\mathcal{E}_{0}$ where the process $W_{t}$ remains finite, but possibly not uniformly bounded. This further implies that on $\mathcal{E}_{0}$, $|S_{t}|/V_{t}\to 0$ and $\limsup_{t}|S_{t}|/\sqrt{V_{t}\ln\ln V_{t}}\leq\sqrt{2}$. In other words, either both of these aforementioned convergences hold or the process $W_{t}$ explodes. In particular, if the data are stochastic with appropriate assumptions, the set $\mathcal{E}_{0}$ is a set of measure $1$, and we conclude that the process $W_{t}$ acts as a witness for the self-normalized SLLN and (upper) LIL.
@@
$\mathcal{E}_{0}:=\left\{V_{t}\uparrow\infty;~\limsup_{t}W_{t}<\infty\right\}.$
@@
path-wise on $\mathcal{E}_{0}$. As a corollary, if the data are drawn from a distribution such that $(S_{t})_{t\geq 1}$ is a sub-Gaussian process with variance $(V_{t})_{t\geq 1}$, then $W_{t}$ is a nonnegative super-martingale, immediately implying that $\mathcal{E}_{0}$ is a measure 1 event, thus implying the self-normalized SLLN and LIL.
As in Theorem 3.3, Theorem 4.2 demonstrates a deterministic statement of the form: either $R_{t}=O(\ln\ln V_{t})$, or the wealth process $W_{t}$ explodes. In case of stochastic data, it provides an explicit super-martingale witnessing the violation of the SLLN and LIL: if either of these is violated on some path, $\limsup_{t}W_{t}=\infty$ on that path.
@@
Since $R_{t}=\frac{S^{2}_{t}}{2V_{t}}-\ln W_{t}$, from (7) we get a lower bound on $W_{t}$. Further, using $\ln(x)\leq x-1$ and completing the squares, we get the following path-wise lower bound on $W_{t}$:
$\ln W_{t}\geq\begin{cases}\frac{S_{t}^{2}}{2V_{t}}-\frac{1}{2}-\ln\tfrac{2}{\ln\ln c}-\ln\ln(c\sqrt{1+V_{t}})-2\ln\ln\ln(c\sqrt{1+V_{t}}),&\text{ if }\frac{|S_{t}|}{V_{t}}\leq\frac{1}{\sqrt{1+V_{t}}}\\ \frac{1}{2}\left(\!\frac{|S_{t}|}{\sqrt{V_{t}}}-\sqrt{1+\frac{1}{V_{t}}}\right)^{2}-\ln\tfrac{2}{\ln\ln c}-\frac{1}{2V_{t}}\\ \qquad\qquad-\ln\ln\left(c\sqrt{1+V_{t}}\right)-2\ln\ln\ln(c\sqrt{1+V_{t}}),&\text{ if }\frac{1}{\sqrt{1+V_{t}}}<\frac{|S_{t}|}{V_{t}}\leq 1\\ \frac{S_{t}^{2}}{2V_{t}}-\frac{V_{t}}{2}\left(\frac{|S_{t}|}{V_{t}}-1+\rho\right)^{2}+\ln\left(\frac{\rho\ln\ln c}{\ln({c}/{1-\rho})(\ln\ln({c}/{1-\rho}))^{2}}\right),&\text{ if }1<\frac{|S_{t}|}{V_{t}}.\end{cases}$ (13)
@@
$\ln W_{t}\geq V_{t}\left(\frac{1-\rho^{2}}{2}\right)+\ln\left(\frac{\rho\ln\ln c}{\left(\ln\tfrac{c}{1-\rho}\right)\left(\ln\ln\tfrac{c}{1-\rho}\right)^{2}}\right).$
Recall that on $\mathcal{E}_{0}$, $V_{t}\uparrow\infty$ and $\limsup_{t}W_{t}<\infty$. From this and the inequality above it, we see that on $\mathcal{E}_{0}$, if $\limsup_{t}\frac{|S_{t}|}{V_{t}}>1$ (i.e., the third branch occurs infinitely often), then $\limsup_{t}\ln W_{t}=\infty$, which leads to a contradiction.
@@
Recall from Section 2.1 that $R_{t}:=f_{t}(\eta^{*}_{t})-\ln W_{t}$, and that for $t$ such that $V_{t}=0$, we also have $S_{t}=0$, and hence for such $t$, for all $\eta\in\mathbb{R}$, $f_{t}(\eta)=0$, and $\ln W_{t}=0$. Thus, the inequalities in the theorem hold trivially in this case. Henceforth, we focus on $t$ so that $V_{t}>0$.